# Irreversibility and symmetry breaking in the creation and annihilation of defects in active living matter

### v1

@@ Abstract
Active living matter continuously creates and annihilates topological defects in a process that remains poorly understood. Here, we investigate these dynamics in two distinct active living systems—swarming bacteria and human bronchial epithelial cells. Despite their entirely different evolutionary origins, biological functions, and physical scales, both systems exhibit half-integer defects, consistent with the nematic phase. However, in contrast to active nematic theory, we find that defect creation and annihilation undergoes spatial symmetry breaking. We propose that this results from a fundamental dualism between nematic structural organization and generated polar forces, which are intrinsic to living systems. Furthermore, our estimation of entropy production reveals that creation and annihilation are not reversed processes. Our findings challenge conventional nematic models and emphasize the role of defect-mediated dynamics in non-equilibrium biological systems as a major source of entropy production.
@@ I Introduction
Active nematic materials are typically comprised of elongated components that generate stress at the microscopic scale, leading to large-scale motion Sanchez et al. (2012); Guillamat et al. (2018); Tan et al. (2019); Kumar et al. (2018). Unlike passive systems, active matter remains in a state of perpetual motion and reorganization, which is driven by internally generated forces Sokolov et al. (2010); Beta et al. (2023). A hallmark of active nematic materials is the continuous creation and annihilation of topological defects Marchetti et al. (2013), which result from a balance between active and elastic forces Giomi (2015); Zhou et al. (2014); Serra et al. (2023); Guillamat Bassedas et al. (2017). Of particular interest are living systems, such as bacterial collectives Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017) and tissue cells Saw et al. (2017); Balasubramaniam et al. (2021); Blanch-Mercader et al. (2018), because their physical properties are also related to their biological functions Be’er and Ariel (2019); Aranson (2022); Doostmohammadi and Ladoux (2022). In living matter, active stress originates from metabolism and the consumption of chemical fuel, which enables bacteria and cells to move and generate forces on their neighbors. Elasticity arises from an elongated cellular shape, leading to liquid-crystal-like properties Be’er and Ariel (2019); Aranson (2022); Doostmohammadi and Ladoux (2022). The resulting defects are mesoscopic objects involving tens to hundreds of cells and governed by intrinsic spatiotemporal scales that are independent of system size Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017); Saw et al. (2017); Balasubramaniam et al. (2021); Blanch-Mercader et al. (2018).
@@
Experiments with bacterial colonies and cell monolayers consistently reveal half-integer defects, in accordance with nematic symmetry Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017); Duclos et al. (2017); Blanch-Mercader et al. (2018); Saw et al. (2017); Kawaguchi et al. (2017); Balasubramaniam et al. (2021); Ienaga et al. (2023). Recent studies suggested that nematic symmetry can also manifest in systems with directionality Venkatesh et al. (2025), such as models of self-propelled rods Shi and Chaté (2018); Großmann et al. (2020) and bacterial experiments Head et al. (2024). Intriguingly, some nematically ordered systems display polar characteristics under specific conditions, indicating the coexistence of nematic and polar symmetries Huber et al. (2018); Meacock et al. (2021); Maroudas-Sacks et al. (2021); Amiri et al. (2022); Lacroix et al. (2024). This interplay of symmetries is fundamental to understanding biological processes, such as biofilm formation and the organization of cellular tissues, yet remains underexplored experimentally.
@@
Defect interactions have been primarily investigated through numerical simulations Thampi et al. (2014); Shankar et al. (2018); Patelli et al. (2019); Bonn et al. (2022); Pearce et al. (2021). Elastic forces between defects predict Coulomb-like interactions, as well as additional forces and torques that rely on their relative orientation Shankar et al. (2022); Pearce and Kruse (2021); Tang and Selinger (2017); Vromans and Giomi (2016). It has been suggested that nematic active stresses also lead to the generation of hydrodynamic flows around the defects which results in self-propulsion of $^{1}/_{2}$ defects. Rotational diffusion can reduce persistence in $^{1}/_{2}$ defect motion, preventing unbinding from $^{1}/_{2}$ counterparts, and affecting defect density in steady states Shankar et al. (2018). However, much of the theoretical framework remains experimentally unverified.
@@
Recent experiments on microtubule-based active nematics and swarming bacteria have uncovered a broad spectrum of defect organization, ranging from giant number fluctuations to hyperuniformity Yashunsky et al. (2024); Fernandez-Nieves et al. (2024); Pearce et al. (2021). In fibrosarcoma cell cultures, $^{1}/_{2}$ defects align toward the edge of the colony, generating chiral edge flows through local chiral active stress Yashunsky et al. (2022). Spontaneous mirror (chiral) symmetry breaking has been demonstrated for $^{1}/_{2}$ defects, which specifically concentrate at the boundary between vorticity-dominated and strain-rate-dominated regions in microtubule-kinesin films, highlighting their strong coupling to large-scale hydrodynamic fluxes within the material Head et al. (2024). However, the creation and annihilation processes, which we identified below as primary sources of energy injection and dissipation in living systems, are underexplored and necessitating experimental validation of current theoretical predictions.
@@
In this paper, we experimentally study swarming bacteria Be’er and Ariel (2019) and human bronchial epithelial cells (HBECs) Blanch-Mercader et al. (2018)—distinct systems that both display nematic half-integer defects. Our findings reveal that the creation and annihilation of defects follow distinct, previously unrecognized trajectories that spontaneously break mirror symmetry—a feature absent in standard active nematic models. We also find that these trajectories exhibit time-reversal symmetry breaking, a hallmark of non-equilibrium processes. Statistical analysis reveals that defect creation and annihilation drive substantial entropy production, underscoring their irreversible nature and critical role in active matter organization. Finally, we demonstrate why these observations can be explained by taking into account directional (polar) self-propulsion in bacteria and cells.
@@ II Results
We study the creation and annihilation of defects in two highly active, quasi-turbulent living systems of swarming Bacillus subtilis bacteria Yashunsky et al. (2024) and human bronchial epithelial cells (HBECs) Blanch-Mercader et al. (2018). Despite the extreme disparity in the time scales and dimension (Fig. 1), both systems share fundamental characteristics of active extensile nematics Giomi (2015); Mart´ınez-Prat et al. (2021): (i) A steady-state mixture of half-integer defects (Figs. 1a-b), which is obtained through continuing creation and annihilation events (Figs. 1c-d), (ii) a counter-rotating vortex pattern around defects with (iii) an exponential distribution of vortex sizes Yashunsky et al. (2022); Blanch-Mercader et al. (2018).
The defects are created and annihilated in pairs with opposite charges. In both systems, the rates of creation and annihilation events fluctuate over time but stabilizes to a steady state (Figs. 1c-d), implying an approximately constant number of defects over time with a total charge that fluctuates around zero. We characterize the motion of defect pairs through these observables, see the sketch in Fig. 2.
@@
  • $\varphi^{+}$: the angle between the orientation of the “tail” of the comet-shaped $^{1}/_{2}$ defect and the vector connecting the pair. Thus, $\varphi^{+}=0$ indicates that the tail is pointing away from the $^{1}/_{2}$ defect.
@@
  • $\varphi^{-}$: the angle between the orientation of one of the $^{1}/_{2}$ legs and the line connecting the pair. We use the smallest angle out of the three legs. For example, $\varphi^{-}=0$ implies that one of the legs is pointing towards the $^{1}/_{2}$ defect.
@@
  • $\delta=\varphi^{+}-3\varphi^{-}-\pi$: pair orientation offset. This gives a measure of how compatible the orientations of the defects are with each other. $\delta=0$ suggests the defects are at the orientations that minimize the variation in the nematic field between themPearce and Kruse (2021).
@@
Figure 2: Configuration of $^{1}/_{2}$ and $^{1}/_{2}$ defect-pairs during creation and annihilation. Sketch: Definitions of orientational angles $\varphi^{+}$ (pink) and $\varphi^{-}$ (violet) for paired $^{1}/_{2}$ and $^{1}/_{2}$ defects relative to the pair axis connecting between $^{1}/_{2}$ and $^{1}/_{2}$ defect centers (horizontal black line). (a, b) Distribution of the pair orientation offset angle, $\delta=\varphi^{+}-3\varphi^{-}-\pi$, as a function of time for creation (blue) and annihilation (red) events. $\delta=0$ implies that the nematic director fields of the paired-defects are synchronized. (c, d) Distribution of $\varphi^{+}$ as a function of time for the creation (blue) and annihilation (red) events. The creation/annihilation instant is indicated by $t=0$, with negative time indicating time to annihilation, and positive time is from creation. Red and blue color intensities indicate probability density. (e-g) Alignment of defect pairs, classified by the mean orientation of the $^{1}/_{2}$ defect during its trajectory, $\langle\varphi^{+}\rangle$. Angular distributions of $\varphi^{+}(d<r_{0})$ and $\varphi^{-}(d<r_{0})$ within $r_{0}=15~\mu m$ for bacteria and $35~\mu m$ for cells are shown for “up” ($\langle\varphi^{+}\rangle>0$, green) and “down” ($\langle\varphi^{+}\rangle<0$, yellow) configurations in bacteria (e, f) and cells (g, h).
@@
Spontaneous breaking of the mirror symmetry in defect orientation
At annihilation or creation events, defect pairs are in phase ($\delta=0$; see Figs. 2a-b), indicating they are perfectly aligned. When $\delta~\neq 0$, the elastic energy between the defects diverges at small separations, indicating that the defects align at very short distances Pearce and Kruse (2021).
@@
We next look at the alignment of the $^{1}/_{2}$ defect with the line connecting the defects in the pair, captured by the parameter $\varphi^{+}$. We see a bimodal distribution for $\varphi^{+}$ immediately before annihilation with peaks at $\pm\pi/2$; see Figs. 2c-d. This indicates that the tail of the positive defect is perpendicular to the direction of annihilation at the end of this process. A similar but less pronounced effect is also visible immediately after creation. This bimodality indicates a spontaneously broken chiral symmetry in the orientation of the defects. To quantify this effect, we split all trajectories into two groups based on their time-averaged orientation across the full trajectory, either “up” (average $\varphi^{+}>0$) or “down” (average $\varphi^{+}<0$). The orientation $\varphi^{-}$ follows a similar symmetry breaking, consistent with $\delta=0$ (Figs. 2e-f).
@@
Figure 3: Spiraling trajectories of creation and annihilation. Sketch: Co-rotation angle $\gamma$—the line connecting between the location of defect centers with respect to the angle close to creation/annihilation instant ($\gamma_{d\rightarrow 0}$). (a, b) Distance between $\pm^{1}/_{2}$ defect-pairs as a function of time after creation (blue) and before annihilation (red), for bacteria and cells. The creation/annihilation instant is indicated by $t({d\rightarrow 0})$, with negative values indicating time to annihilation, and positive ones indicating time from creation. Solid lines represent the average trajectories, while red and blue color intensities indicate probability density. (c, d) Time evolution of $\gamma$ (the angle of the line connecting the defect-pairs, aligned to $\gamma_{d\rightarrow 0}=0$). Solid lines represent averages for “up” (green) and “down” (yellow) configurations. Shaded areas indicate the standard deviation. (e, f) Average trajectories of $^{1}/_{2}$ (pink) and $^{1}/_{2}$ (violet) defects during creation and annihilation, measured in the lab frame of reference. Trajectories are centered around the creation/annihilation position (black point), rotated to $\gamma_{d\rightarrow 0}$ and mirrored to the “up” configuration.
@@
The separation rate between paired defects (Figs. 3a-b) is characterized by increased speed close to the creation or annihilation events, consistent with theoretical predictions for defect annihilation in passive systems Pearce and Kruse (2021). However, defects do not move in a straight line. This is quantified by the orientation of the vector connecting the defects changes over time, $\gamma(t)$. Generally, $\gamma$ varies monotonically with time, indicating that defect pairs tend to co-rotate during the process of annihilation or creation. Furthermore, the chirality of the co-rotation is correlated with the “up” or “down” configuration (Figs. 3c-d).
Lastly, Figs. 3e-f show averaged trajectories for creation and annihilation events in the “up” configuration. This reflects the spiraling motion toward (away) each other during annihilation (creation). These trajectories are similar to those predicted for the annihilation of topological defects with an initial phase disparity Pearce and Kruse (2021).
Figure 4: Ensemble-averaged flow fields near $\pm^{1}/_{2}$ defect pairs. (a, b) Average velocity fields (black arrows) and vorticity distributions (heatmap) around defects. (a-d) Flow fields around paired $^{1}/_{2}$ and $^{1}/_{2}$ defects, averaged separately for the “up” (a, b) and “down” (a, d) configurations. e, f) Flow fields around unpaired $^{1}/_{2}$ and $^{1}/_{2}$ defects. For clarity, only results for bacteria close to annihilation are shown. The angular orientation of defects is illustrated in the schematic. Arrow and spatial scales are indicated. Scale bar: $10~\mu m$. The number of fields averaged to obtain each panel is: (a) $5459$; (b) $5391$; (c) $5336$; (d) $5323$; (e) $89027$; (f) $88267$.
Spontaneous breaking of the mirror symmetry in flow fields
We find that flows around defect pairs also exhibit mirror symmetry breaking (Figs. 4a-d, Extended Data Figs. 1-2). Flows were measured separately for the “up” and “down” orientations. In bacteria, the forward flow at the core of the defect $^{1}/_{2}$ is offset relative to the defect symmetry axis. The direction of the tilt changes between annihilation and creation (Extended Data Fig. 1). In particular, for both bacteria and cells, the flow at the core of the $^{1}/_{2}$ defects exhibits a distinct rotational behavior, rotating clockwise for “up”-oriented configurations (Fig. 4B, Extended Data Figs. 1b,f and 2b,f) and counterclockwise for “down”-oriented configurations (Fig. 4d, Extended Data Figs. 1d,h and 2d,h). The flows around unpaired defects (Figs. 4e-f) retain mirror symmetry, in agreement with theoretically predicted flows around isolated defects in extensile active nematics.
@@
Figure 5: The interaction between nematic and polar symmetries. (a) Defect creation via bend instability with the overlaid polar field, indicated by red arrows; a grain boundary in the polar field is introduced indicated by the blue dashed line. The polar field leads to a torque on the defects, indicated by the green arrow. The torque will reorient the $^{1}/_{2}$ defect, causing it to self-propel around the $^{1}/_{2}$ defect. Shown here is the “up” configuration. (b) Nematic and polar fields around a pair of defects before annihilation. Here the grain boundaries do not directly connect. The same torque is present around the $^{1}/_{2}$ defect that can lead to a spiral trajectory. (c) & (d) Predicted Stokes flow field around a (c) $^{1}/_{2}$ and (d) $^{1}/_{2}$ defect with additional active force along the polar direction; shown in the “up” configuration.
@@
In addition to nematic order, bacteria and cells also exhibit a polar symmetry due to their directed self-propulsion mechanism. Combining nematic interactions with active stresses of various symmetries can significantly alter active nematic behavior Ardaševa et al. (2025); Großmann et al. (2020).

### v2

@@ Abstract
Active living matter continuously creates and annihilates topological defects in a process that remains poorly understood. Here, we investigate these dynamics in two distinct active living systems—swarming bacteria and human bronchial epithelial cells. Despite their entirely different evolutionary origins, biological functions, and physical scales, both systems exhibit half-integer defects, consistent with the nematic phase. However, in contrast to active nematic theory, we find that defect creation and annihilation undergoes spatial symmetry breaking. We propose that this results stems from a fundamental dualism between nematic structural organization and generated polar forces, which are intrinsic to living systems. Furthermore, estimation of entropy production reveals that creation and annihilation are not reversed processes. Our findings challenge conventional nematic models and emphasize the role of defect-mediated dynamics in non-equilibrium biological systems as a major source of entropy production.
@@ I Introduction
Active nematic materials are typically comprised of elongated components that generate stress at the microscopic scale, leading to large-scale motion Sanchez et al. (2012); Guillamat et al. (2018); Tan et al. (2019); Kumar et al. (2018). Unlike passive systems, active matter remains in a state of perpetual motion and reorganization, which is driven by internally generated forces Sokolov et al. (2010); Beta et al. (2023). A hallmark of active nematic materials is the continuous creation and annihilation of topological defects Marchetti et al. (2013), which result from a balance between active and elastic forces Giomi (2015); Zhou et al. (2014); Serra et al. (2023); Guillamat Bassedas et al. (2017). Of particular interest are living systems, such as bacterial collectives Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017) and tissue cells Saw et al. (2017); Balasubramaniam et al. (2021); Blanch-Mercader et al. (2018), because their physical properties are also related to their biological functions Be’er and Ariel (2019); Aranson (2022); Doostmohammadi and Ladoux (2022). In living matter, active stress originates from metabolism and the consumption of chemical fuel, which enables bacteria and cells to move and generate forces on their neighbors. Elasticity arises from an elongated cellular shape, leading to liquid-crystal-like properties Be’er and Ariel (2019); Aranson (2022); Doostmohammadi and Ladoux (2022). The resulting defects are meso-scopic objects involving tens to hundreds of cells and governed by intrinsic spatiotemporal scales that are independent of system size Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017); Saw et al. (2017); Balasubramaniam et al. (2021); Blanch-Mercader et al. (2018).
@@
Experiments with bacterial colonies and cell monolayers consistently reveal half-integer defects, in accordance with nematic symmetry Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017); Duclos et al. (2017); Blanch-Mercader et al. (2018); Saw et al. (2017); Kawaguchi et al. (2017); Balasubramaniam et al. (2021); Ienaga et al. (2023). Recent studies suggested that nematic symmetry can also manifest in systems with clear polar directionality Venkatesh et al. (2025), such as self-propelled rod models Shi and Chaté (2018); Großmann et al. (2020); de Graaf Sousa et al. (2025) and rod-like bacteria gliding on surfaces Han et al. (2025). Under certain conditions, active nematic systems can even acquire polar features, indicating that nematic and polar symmetries can coexist Huber et al. (2018); Meacock et al. (2021); Maroudas-Sacks et al. (2021); Amiri et al. (2022); Lacroix et al. (2024); Vafa and Doostmohammadi (2025). In Han et al. (2025), polar stresses in the nematically organized Myxococcus xanthus colonies were shown to be associated with multi-layer formation. Understanding how living systems organize may depend on this interplay of symmetries, which could underlie processes such as biofilm formation and tissue organization, yet it remains largely untested experimentally.
@@
Defect interactions have been primarily investigated through numerical simulations Thampi et al. (2014); Shankar et al. (2018); Patelli et al. (2019); Bonn et al. (2022); Pearce et al. (2021). Elastic forces between defects result in polar-like interactions, as well as additional forces and torques that rely on their relative orientation Shankar et al. (2022); Pearce and Kruse (2021); Tang and Selinger (2017); Vromans and Giomi (2016). It has been observed that nematic active stresses also lead to the generation of hydrodynamic flows around the defects which results in self-propulsion of $^{1}/_{2}$ defects. Rotational diffusion can reduce persistence in $^{1}/_{2}$ defect motion, preventing unbinding from $^{1}/_{2}$ counterparts, and affecting defect density in steady states Shankar et al. (2018). However, much of the theoretical framework remains experimentally unverified.
@@
Recent experiments on microtubule-based active nematics and swarming bacteria have uncovered a broad spectrum of defect organization, ranging from giant number fluctuations to hyperuniformity Yashunsky et al. (2024); de la Cotte et al. (2025); Pearce et al. (2021). In fibrosarcoma cell cultures, $^{1}/_{2}$ defects align toward the edge of the colony, generating chiral edge flows through local chiral active stress Yashunsky et al. (2022). Spontaneous mirror (chiral) symmetry breaking has been demonstrated for $^{1}/_{2}$ defects, which specifically concentrate at the boundary between vorticity-dominated and strain-rate-dominated regions in microtubule-kinesin films, highlighting their strong coupling to large-scale hydrodynamic fluxes within the material Head et al. (2024). However, the creation and annihilation processes, which we identify below as primary sources of energy injection and dissipation in living systems, are underexplored, necessitating experimental validation of current theoretical predictions.
@@
In this paper, we experimentally study swarming bacteria Be’er and Ariel (2019) and human bronchial epithelial cells (HBECs) Blanch-Mercader et al. (2018)—two distinct systems that are known to display nematic half-integer defects. Our findings reveal that the creation and annihilation of defects follow distinct, previously unrecognized trajectories that spontaneously break mirror symmetry—a feature absent in standard active nematic models. We also find that these trajectories exhibit time-reversal symmetry breaking, a hallmark of non-equilibrium processes. Statistical analysis reveals that defect creation and annihilation are a major source of entropy production in the bulk, underscoring their irreversible nature and critical role in active matter organization. Finally, we propose a theoretical model that explains these observations taking into account directional (polar) self-propulsion in bacteria and cells.
II Results
@@
We study the creation and annihilation of defects in two highly active, quasi-turbulent living systems of swarming Bacillus subtilis bacteria Yashunsky et al. (2024) and human bronchial epithelial cells (HBECs) Blanch-Mercader et al. (2018). Despite the extreme disparity in the time scales and dimension (Fig. 1), both systems share fundamental characteristics of active extensile nematics Giomi (2015); Martínez-Prat et al. (2021): (i) A steady-state mixture of half-integer defects (Figs. 1a-b), which is obtained through continuing creation and annihilation events (Figs. 1c-d), (ii) a counter-rotating vortex pattern around defects with (iii) an exponential distribution of vortex sizes Yashunsky et al. (2022); Blanch-Mercader et al. (2018).
The defects are created and annihilated in pairs with opposite charges. In both systems, the rates of creation and annihilation events fluctuate over time but stabilizes to a steady state (Figs. 1c-d), implying an approximately constant number of defects over time with a total charge that fluctuates around zero. We characterize the motion of defect pairs through these observables, see the sketches in Figs. 2 and 3.
@@
  • $\varphi^{+}$: the angle between the orientation of the “tail” of the comet-shaped $^{1}/_{2}$ defect and the vector connecting the pair. Thus, $\varphi^{+}=0$ indicates that the tail is pointing toward the $^{1}/_{2}$ defect.
@@
  • $\varphi^{-}$: the angle between the orientation of one of the $^{1}/_{2}$ legs and the vector connecting the pair. We use the smallest angle out of the three legs. For example, $\varphi^{-}=0$ implies that one of the legs is pointing away from the $^{1}/_{2}$ defect.
@@
  • $\delta=\varphi^{+}-3\varphi^{-}-\pi$: pair orientation offset. This gives a measure of how compatible the orientations of the defects are with each other. $\delta=0$ suggests the defects are at the orientations that minimize the variation in the nematic field between them Pearce and Kruse (2021); Tang and Selinger (2017); Vromans and Giomi (2016).
@@
Figure 2: Configuration of $^{1}/_{2}$ and $^{1}/_{2}$ defect-pairs during creation and annihilation. Center sketch: Definitions of orientational angles $\varphi^{+}$ (pink) and $\varphi^{-}$ (violet) for paired $^{1}/_{2}$ and $^{1}/_{2}$ defects. These angles are defined relative to the vector pointing from the $^{1}/_{2}$ to the $^{1}/_{2}$ defect, shown by the solid black arrow. (a, b) Distribution of the pair orientation offset angle, $\delta=\varphi^{+}-3\varphi^{-}-\pi$, as a function of time for creation (blue) and annihilation (red) events. $\delta=0$ implies that the nematic director fields of the paired-defects are in phase. (c, d) Distribution of $\varphi^{+}$ as a function of time for the creation (blue) and annihilation (red) events. The creation/annihilation instant is indicated by $t=0$, with negative time indicating time to annihilation, and positive time is from creation. Red and blue color intensities indicate probability density. The bright line close to $t=0$ indicates a region in which the distance between the defects is up to half cell-body length ($3-4~\mu m$ for bacteria and $7-8~\mu m$ for cells). At these distances the continuous nematic field is not well defined and the discrete nature of particles is of concern. (e-g) Angular distribution of $\varphi^{+}$ and $\varphi^{-}$ for defect pairs at small separation ($d<15~\mu m$ for bacteria and $d<35~\mu m$ for cells). Pairs are classified as “up” ($\langle\varphi^{+}\rangle>0$, green) or “down” ($\langle\varphi^{+}\rangle<0$, yellow).
@@
Spontaneous breaking of mirror symmetry during defect creation and annihilation
@@
At the point of annihilation or creation, defect pairs are in phase ($\delta=0$; see Figs. 2a-b), indicating they are perfectly aligned. This is due to elastic torques that diverge when defects are out of phase at small distances Pearce and Kruse (2021). However, the orientation of the $^{1}/_{2}$ defect relative to the position of the $^{1}/_{2}$ defect (given by $\varphi^{+}$) is free to vary, provided they remain in phase. In a typical active nematic system, the average orientation of the $^{1}/_{2}$ defect during creation or annihilation is dictated by the active stress, which causes the $^{1}/_{2}$ defect to self-propel; for an extensile active nematic one expects $\varphi^{+}=0$ during creation and $\varphi^{+}=\pm\pi$ during annihilation Giomi et al. (2014); Shankar et al. (2018). Conversely, in both our bacteria and cellular experiments, we observe a bimodal distribution of $\varphi^{+}$ immediately before annihilation with peaks around $\pm\pi/2$; see Figs. 2c-d. This indicates that the tail of the positive defect is oriented roughly perpendicular relative to the direction of annihilation at the end of this process. Since $\delta=0$, $\varphi^{-}$ varies by about one-third as much. A similar but less pronounced effect is also visible immediately after creation. This bimodality indicates a spontaneously broken mirror symmetry in the process of creation and annihilation. To quantify this effect, we split all trajectories into two groups based on their time-averaged orientation across the full trajectory, either “up” ($\langle\varphi^{+}\rangle>0$) or “down” ($\langle\varphi^{+}\rangle<0$). The orientation $\varphi^{-}$ follows a similar symmetry breaking, consistent with $\delta=0$ (Figs. 2e-g).
In contrast, active tubulin-kinesin suspensions, which are purely nematic, exhibit no such mirror-symmetry breaking behavior. Analysis of data from Tan et al. Tan et al. (2019) (Extended Figs. 4c and 5) shows that the distribution of $\varphi^{+}$ has a single peak around 0 (creation) or $\pi$ (annihilation), in accordance with the prediction of active nematic theory.
@@
Figure 3: Spiraling trajectories of creation and annihilation. Center sketch: Co-rotation angle $\gamma$—the line connecting between the location of defect centers with respect to the angle close to the creation/annihilation instant ($\gamma_{d\rightarrow 0}$). (a, b) Distance between $\pm^{1}/_{2}$ defect-pairs as a function of time after creation (blue) and before annihilation (red), for bacteria and cells. The creation/annihilation instant is indicated by $t({d\rightarrow 0})$, with negative values indicating time to annihilation, and positive ones indicating time from creation. Solid lines represent the average trajectories, while red and blue color intensities indicate probability density. (c, d) Time evolution of $\gamma$ (the angle of the line connecting the defect-pairs, aligned to $\gamma_{d\rightarrow 0}=0$). Solid lines represent averages for “up” (green) and “down” (yellow) configurations. Shaded areas indicate the standard deviation. (e, f) Average trajectories of $^{1}/_{2}$ (pink) and $^{1}/_{2}$ (violet) defects during creation and annihilation, measured in the lab frame of reference. Trajectories are centered around the creation/annihilation position (black point), rotated to $\gamma_{d\rightarrow 0}=0$ and mirrored to the “down” configuration.
Figure 4: Symmetry breaking in the flow around $^{1}/_{2}$. (a–b) The distribution of $\phi_{V}$ in bacteria, the angle between the $^{1}/_{2}$ velocity and defect orientation (inset in panel a). Average flow fields around $^{1}/_{2}$ defects are obtained by separately averaging trajectories according to their defect-pair configurations: “up” ($\langle\varphi^{+}\rangle>0$) or “down” ($\langle\varphi^{+}\rangle<0$). Flows are shown for defects with overall displacement to the right (blue–gray) and to the left (pink). The velocity scale indicated by the black arrow is $10~\mu\mathrm{m}/\mathrm{s}$. Scale bar: $10~\mu\mathrm{m}$.
@@
The separation rate between paired defects (Figs. 3a-b) is characterized by increased speed close to the creation or annihilation events, consistent with theoretical predictions for defect annihilation Giomi et al. (2014). However, contrary to typical active nematics, defects do not move in a straight line, but rather spiral toward or away from each other, see Figs. 3e-f for average trajectories of defects in the “down” configuration. The handedness of the spiral motion can be quantified by plotting $\gamma(t)$; the orientation of the vector connecting the defects. Generally, $\gamma$ varies monotonically with time, indicating that defect pairs tend to co-rotate during the process of annihilation or creation. Furthermore, the direction of the co-rotation is correlated with the “up” or “down” configuration (Figs. 3c-d). These spiral paths are similar to those predicted for the annihilation of topological defects with an initial phase disparity Pearce and Kruse (2021); Vromans and Giomi (2016), however since we do not observe any major phase disparity (Figs. 2a-b) we predict that the spiraling arises from active forces and self-propulsion of the $^{1}/_{2}$ defect.
Spontaneous breaking of mirror symmetry in the flow fields
self-propulsion of $^{1}/_{2}$ defects arises from the characteristic flow fields generated around the defect by the active forces. For a typical active nematic, the flow around a $^{1}/_{2}$ defect features two counter-rotating vortices and a central flux aligned with the defect which causes it to self-propel Giomi et al. (2014).
In order to quantify the direction in which defect move we introduce two additional angles to be measured.
  • $\phi_{V}$: For a single $^{1}/_{2}$ defect, $\phi_{V}$ is the angle between the defect velocity (calculated from the displacement between consecutive frames) and the defect orientation.
In words, $\phi_{V}$ is the offset between the orientation of a $^{1}/_{2}$ defect and the direction in which it moves. The theory of active nematics predicts that for an isolated defect $\phi_{V}=\pm\pi$ Giomi et al. (2014).
Figure 4 shows the direction of the $^{1}/_{2}$ movement, $\phi_{V}$, during creation and annihilation events. The distribution of angles is bimodal, suggesting that the spontaneous break in symmetry is also manifested in a slanted trajectory (compared to the defect orientation). This can be directly observed by averaging the flows around recently created/soon to be annihilated $^{1}/_{2}$’s, partitioned according to their “up” or “down” configurations (Fig. 4).
We find low, yet highly significant correlation between $\phi_{V}$ close to creation and close to annihilation of the same defect $^{1}/_{2}$ (for details see Extended Data Fig. 6). This indicates that the direction off-axis motion is consistent throughout the lifetime of the defect.
Overall, the results above suggest that $^{1}/_{2}$ are created with an “up” or “down” orientation. This orientation persists throughout the life of the defect and finally determines its configuration during annihilation. The configuration of the $^{1}/_{2}$ seems to be constrained or enslaved to the $^{1}/_{2}$, Extended Data Figs. 1 and 2
Finally, comparing with active nematic tubulin-kinesin suspensions, which do not exhibit mirror-symmetry breaking, the displacement of $^{1}/_{2}$ defects remains aligned with their orientation. This is evidenced by the single peak in the distribution of $\phi_{V}$ at $\pi$ (Extended Fig. 6). This demonstrates that the mirror-symmetry breaking observed in bacteria and cells does not arise from the presence of counter-rotating vortex structures, which are generic to active nematic turbulence Doostmohammadi et al. (2016).
@@
Figure 5: The interaction between nematic and polar symmetries. (a) Defect creation via bend instability with the overlaid nematic (black lines) and polar (red arrows) fields. A grain boundary in the polar field is introduced (blue dashed line). The polar field leads to a torque on the defects (green arrow), which will reorient the $^{1}/_{2}$ defect, causing it to self-propel and spiral around the $^{1}/_{2}$ defect (inset). Shown here is the “down” configuration. $\zeta_{p}\neq 0$. (b) Nematic and polar fields around a pair of defects before annihilation. Here, the grain boundaries do not directly connect. A similar torque, acting on the $^{1}/_{2}$ defect, leads to a spiraling trajectory (inset). (c) Total co-rotation ($\Delta\gamma$) and (d) final positive defect orientation ($\Delta\varphi^{+}$) of the defect pair as a function of the applied polar body force ($\zeta_{p}$). Dashed lines indicate creation and solid lines for annihilation, green “up” and yellow “down” configuration. (e) Predicted Stokes flow around a $^{1}/_{2}$ defect in the annihilation “up” (or creation “down”) configuration. Shown with $|\zeta_{p}/\zeta_{Q}|=0.5$ and $\phi_{g}=0$. (f) Angular offset of defect self-propulsion ($\phi_{V}$) and (g) net vorticity around a $^{1}/_{2}$ defect as a function of the relative polar body force ($\zeta_{p}/\zeta_{Q}$) and position of the grain boundary ($\phi_{g}$). $\phi_{g}=0$ implies that the grain boundary alignment with the tail of the $^{1}/_{2}$ defect is as in (a) and (b).